# From Acceleration to Saturation: Scaling Behavior of Bootstrapped Language Model Pretraining

### v1

@@ 3 Formulating the Data Scaling Laws
Our goal is to find a functional form $L(D_{1},D_{2})$ that captures how the second-stage loss depends jointly on both stages. To derive a principled formulation, we impose two natural constraints:
@@
$L(D_{1},D_{2})=L_{D_{1}}(D_{2})=AD_{2}^{-\alpha}+E.$ (1)
This is consistent with classical neural scaling laws and reflects the expectation that, when starting from a fixed initialization, additional tokens improve performance predictably.
Condition 2: For any fixed value of $D_{2}$, the loss should exhibit power-law behavior with respect to the number of first-stage tokens $D_{1}$: $L(D_{1},D_{2})=L_{D_{2}}(D_{1})=AD_{1}^{-\alpha}+E$. This is consistent with the power-law ansatz and captures the intuition that a better-trained base model (i.e., larger $D_{1}$) should result in lower loss. Moreover, this condition implies that as $D_{2}\to 0$, the loss should continuously approach that of the base model: $\lim_{D_{2}\to 0}L(D_{1},D_{2})=AD_{1}^{-\alpha}+E$, which is a natural requirement for:  [^2] As in Footnote 1, we omit the +1+1 term in $D+1$ expressions for notational simplicity, such that terms proportional to $D_{2}$ are finite as $D_{2}\to 0$.
  • CPT: The second stage’s initial loss should begin from the base model’s (evaluated on second-stage dataset).
  • Function-preserving model growth: When model capacity is expanded but initialized carefully, the initial loss should remain close to the base model’s loss.
@@
$\textbf{Multiplicative:}\quad L(D_{1},D_{2})$ $=AD_{1}^{-\alpha_{1}}D_{2}^{-\alpha_{2}+\alpha_{3}\log D_{1}}+E.$ (2)
$\textbf{Additive:}\quad L(D_{1},D_{2})$ $=AD_{1}^{-\alpha_{1}}+FD_{2}^{-\alpha_{2}}+E.$ (3)
$\textbf{Hybrid:}\quad L(D_{1},D_{2})$ $=\big(AD_{1}^{-\alpha_{1}}+F\big)D_{2}^{-\alpha_{2}}+E.$ (4)
@@
Note that since $D_{1}^{-\alpha_{1}}D_{2}^{-\alpha_{2}+\alpha_{3}\log D_{1}}=D_{1}^{-\alpha_{2}+\alpha_{3}\log D_{2}}D_{2}^{-\alpha_{2}}$, the multiplicative form with an interaction term satisfies our conditions, while other forms do not. We further note that multiplicative scaling laws have been observed in mikami2022scaling ; zhang2024scaling but without the interaction term. liewscaling empirically showed that sparse upcycling (training sparse mixture-of-experts (MoE) models reusing existing dense models) follows a scaling law similar to Equation 2 but under different motivation and conditions. Our work directly extends their findings to other pretraining paradigms. Equation 4 has also been studied in barnett2024empirical .
4 Data Scaling Laws
Fitting the scaling laws. To determine the functional form described in the previous Section that best describes the scaling behavior of bootstrapped pretraining methods, we use a model of size 0.1B, trained on a $5\times 5$ grid of $D_{1},D_{2}$. For CPT, we consider both code and mathematics datasets, while for model growth, we consider both width expansion and depth-wise stacking that double the size of the base model.  [^3] We increase the size of the hidden dimension by $\sqrt{2}$ for width expansion; and we double the number of non-embedding layers for stacking.
Running the experiments as described above and obtaining the results, we fit the losses with the functional forms of Equations 2, 3, and 4. Following hoffmann2022training ; besiroglu2024chinchilla , we perform optimization using the Huber loss ($\delta=10^{-3}$) and the BFGS algorithm, to fit the logarithm of the loss via the LogSumExp trick applied to the RHS of functional forms. The leave-one-out root mean square error (RMS) serves as the goodness-of-fit metric.
As can be seen in Table 1, the multiplicative form with interaction, Equation 2, consistently achieves the lowest error across all methods and datasets, indicating that it best captures the underlying scaling behavior.
Bottom-up empirical evidence. To strengthen our proposal for Equation 2, we demonstrate through a bottom-up observational approach (instead of the top-down approach of deriving the scaling laws by imposing suitable conditions) that the multiplicative nature of the scaling law, including the interaction term, arises naturally from empirical data.
In Figure 1, we first perform power-law fits to show that the loss follows Equation 1 for fixed $D_{1}$, validating Condition 1 we impose for deriving the scaling laws. More importantly, we observe that the fitted scaling exponent of Equation 1 decreases as $D_{1}$ increases, with the model growth method more pronouncedly so. To quantify this trend, we express (the minus of) the scaling exponent as a function of $D_{1}$, denoted by $\alpha(D_{1})$. A scatter plot of $\alpha(D_{1})$ reveals a clear logarithmic dependence:
$\alpha(D_{1})=\gamma\log D_{1}+E^{\prime}.$ (5)
This relationship fits the empirical data well, as shown in the same Figure. Substituting Equation 5 into Equation 1, we arrive at a term of the form $D_{2}^{-E^{\prime}+\gamma\log D_{1}}$, providing direct empirical justification for the interaction term. This demonstrates that the interaction term is not a theoretical artifact, but rather a necessary component for capturing observed trends in the data. Additionally in the same Figure, we show that the multiplicative dependence, $A\propto D_{1}^{-\alpha_{1}}$ also arises from empirical observations.
We further show that the loss follows a power law with respect to $D_{1}$ for fixed $D_{2}$ in Figure 3 (with more plots in Appendix D). This validates Condition 2 we impose for deriving the scaling laws.
Figure 3: $D_{1}$ has power-law scaling. We show scaling behavior of first-stage training tokens ($D_{1}$) for different values of second-stage tokens ($D_{2}$), indicating that $D_{1}$ also has power-law scaling. Left: Continual pretraining on code data; Right: model growth by stacking. More plots in Appendix D.
Table 1: Multiplicative scaling law with interaction consistently achieves lowest error. Leave-one-out RMS error ($\times 10^{-3}$) for fitting the loss for reusing a 0.1B model. Functional forms are from Equations 2, 3, and specific cases with $\alpha_{3}=0$.
RMS $(\times 10^{-3})$ Mul. Mul. ($\alpha_{3}=0$) Add. Hyb.
CPT (code) 1.573 2.095 3.913 2.245
CPT (math) 1.737 3.147 7.443 3.340
Expand 2.790 4.837 9.855 4.841
Stack 2.845 7.818 10.323 7.748
4.1 Interpretations
Several quantitative and qualitative insights can be made from the scaling law. Fixing $D_{1}$, the effective scaling factor for $D_{2}$ becomes $AD_{1}^{-\alpha_{1}}$. As $D_{1}$ increases, this factor decreases, resulting in a lower initial loss at the start of second-stage pretraining, agreeing with the conventional wisdom that better-pretrained base models (larger $D_{1}$) provide stronger initialization for second-stage pretraining.
As shown in Table 7 in Appendix D, the fitted values of $\alpha_{3}$ are consistently positive across all settings and datasets, reinforcing the generality of this saturation phenomenon (see also Section 5.1 for additional evidence). Also note that $\alpha_{3}$ is typically at least an order of magnitude smaller than other exponents (with its effects accumulating logarithmically with $D_{1}$), which may explain why it has not been observed widely in previous scaling studies.
5 A Closer Look At The Scaling Behavior
We first perform a more detailed analysis of the scaling behavior of bootstrapped pretraining methods, by examining more variants of scaling law and training methods. The aim is to validate the robustness of the multiplicative scaling law with interaction. We then propose and validate a joint scaling law of dataset and model sizes. The rest of the Section is subsequently dedicated to studying the practical implications of the scaling laws.
5.1 Variants of Bootstrapped Pretraining
Variant of CPT scaling law. When the dataset used for CPT is the same as the base model, we expect the scaling formula shown below holds:
$L(D_{1},D_{2})=A(D_{1}+D_{2})^{-\alpha}+E$ (6)
Hence, one may expect this holds for CPT on a different dataset as well, as assumed in que2024d ; wang2025learning . Fitting the above formula for the code (math) CPT scenario, we find that the RMS error is 0.0213 (0.0235), higher than the best ones in Table 1. Hence, our scaling law models CPT better, in contrast to previous assumptions, which may have overlooked the overtraining effects on the model.
Variants of CPT. We also consider CPT variants that are commonly used in practice ibrahim2024simple :
@@ 5.1 Variants of Bootstrapped Pretraining
  • CPT with replay: A portion of the first-stage data is mixed into the second-stage data, which is common for mitigating catastrophic forgetting. We consider a replay ratio of 0.25, i.e., 25% of the second-stage data is from the first stage.
@@
  • CPT from stable phase: We consider CPT starting from a base model checkpoint in the stable phase of the WSD learning rate schedule. This is to avoid adverse effects from re-warming the learning rate from a decayed value.
Variants of model growth. We further consider model growth variants.
  • Larger growth factor: We consider a larger growth factor of 4 times (instead of 2).
@@
  • Model growth from stable phase: Similar to CPT, we train the stacked model starting from a base model checkpoint in the stable phase of the WSD learning rate schedule. This also avoids re-warming the learning rate from a decayed value.
We again conduct experiments using a 0.1B model, and fit the scaling laws with all possible functional forms as in Section 4 for the above variants. Results are summarized in Table 2, with fitted scaling exponents also shown in Table 7 of Appendix D. These results show that the propsed scaling law still has the lowest RMS, and variants like replay and larger growth factor do not change the functional form. Furthermore, the saturation effects are unlikely to be artifacts of re-warmings of the learning rate. Overall, the proposed scaling law is robust across datasets, configurations, and methods investigated.
Table 2: Multiplicative scaling law with interaction is robust over various configurations (RMS Error). Leave-one-out RMS error ($\times 10^{-3}$) for fitting the loss for variants of bootstrapped pretraining. See Table 1 for definitions and Section 5 for variant details.
RMS $(\times 10^{-3})$ Mul. Mul. ($\alpha_{3}=0$) Add. Hyb.
CPT (replay) 0.987 2.222 4.646 2.223
CPT (stable) 1.371 2.366 4.315 2.367
Stack (x4) 2.152 6.557 8.866 6.667
Stack (stable) 2.957 8.757 10.559 8.822
@@ 5.2 Joint Scaling Incorporating Model Size
Let us first focus on CPT where the model size remains unchanged after bootstrapping. We proceed to extend the data scaling law to include model size $N$. To determine the functional form, we impose that the loss follows the well-grounded “Chinchilla" scaling law hoffmann2022training (which jointly models the base model’s loss with respect to dataset and model sizes) with respect to $D_{2}$ and $N$:
$L_{D_{1}}(D_{2},N)=AD_{2}^{-\alpha}+N^{-\beta}+E.$
@@
Fitting the joint scaling law. We conduct experiments with base models of sizes 15 million (M), 44M, 0.1B, 0.2B, 0.5B and 1B, training them with different numbers of first and second-stage tokens as in Section 4. To fit the joint scaling law, we use the same fitting procedure as in Section 4, additionally taking $N$ as an additional variable. In the left panel of Figure 4, we show that the formula fits the data well for CPT on code, and the goodness-of-fit to the validation data points also shows that it can be extrapolated to larger model and dataset sizes. See Table 8 in Appendix D.2 for the fitted coefficients. In the same Appendix, we further extend our results to model growth where model sizes increase.
Figure 4: Left: Joint scaling law fit for continual pretraining on code data. Orange points indicate the 10% of data with lowest losses used for validation. Right: Model growth efficiency decreases with sunk cost and model size. For token budgets above the curve(s), training from scratch is more efficient than stacking-based model growth; for budgets below, model growth remains advantageous. Shown are the numerical (blue) and analytical (red) solutions of Equation 8.
5.3 Practical Implications
We consider two practical implications based on the extrapolation of the scaling laws.
Training from scratch versus bootstrapped pretraining. We show that by comparing the scaling laws with the base model scaling laws, one can determine whether bootstrapped pretraining is more efficient than training from scratch under different circumstances.
A key practical consideration is how much of the initial investment in pretraining, the so-called sunk cost ($D_{1}$), can be effectively leveraged in the second stage of training. This question has been explored in the context of MoE upcycling strategies komatsuzaki2022sparse ; liewscaling . For example, it was found that upcycling yields benefits up to 120% of the sunk cost. That is, to match the performance of an upcycled MoE model that underwent an additional 0.4 trillion tokens of training after an initial 2T tokens, training a comparable MoE from scratch would require 2.4T tokens—representing an effective saving of 2T tokens.
To investigate this in model growth, we define $D^{*}$ as the number of tokens required for training from scratch to match the performance of a grown model with the same sunk cost, following liewscaling :
$L^{\rm scratch}_{2N}(D^{*})=L^{\rm grown}_{N}(D_{1}=D^{*},D_{2}=D^{*})$ (8)
For CPT on code data, we can perform similar extrapolation as well, finding that a 7B model trained from scratch begins to outperform bootstrapped models (pretrained on 10T tokens) once more than 300B tokens of domain-specific data are used. See Appendix E for details. These findings offer quantitative guidance for deciding between bootstrapped pretraining and training from scratch (or from intermediate checkpoints), depending on the extent to which the base model is overtrained, as well as the available budget.
Compute optimality of bootstrapped pretraining. For scaling studies of LLMs, the cost of training is often estimated using floating point operations (FLOPs), which we simply refer to as compute. The compute cost can be approximated as $C=6ND$ kaplan2020scaling . One is often interested in training LLMs with the least amount of compute. For CPT, optimizing the compute based on Equation 2 leads to the scaling relations:
where $\alpha_{\rm eff}\coloneq\alpha_{2}-\alpha_{3}\log D_{1}$. Notably, as $D_{1}$ increases, $\alpha_{\rm eff}$ decreases, meaning larger models require more tokens for compute-optimal bootstrapped pretraining. We give a similar derivation for model growth in Appendix E. We leave empirical verification of these relationships to future work, primarily due to cost considerations (see Appendix B.4 for an estimate of GPU hours spent in the work).
6 Discussion and Conclusion
In this paper, we have shown that while bootstrapped pretraining methods can accelerate learning, they exhibit scaling saturation as the base model is overtrained. Our work bears practical implications as even relatively tiny models are overtrained to trillions of tokens nowadays zhang2024tinyllama ; dubey2024llama , and bootstrapped pretraining is widely used in practice. We encourage the community to report the amount of pretraining data used for the base model, and even make intermediate checkpoints available, which can be more useful than the final model for bootstrapped pretraining.
We have provided a fairly broad scaling study of two-stage pretraining and there are many potential future directions worth diving deeper into. Let us highlight a few of them.
Incorporating other factors into the scaling laws. We have not considered the scaling behavior with respect to other method-specific factors, such as the CPT replay ratio que2024d ; wang2025learning and the growth factor in model growth du2024stacking , as (1) our focus is on the more broadly applicable scaling behavior (i.e., scaling with respect to training tokens), and (2) adding more factors would increase the complexity of the scaling laws and the number of parameters to fit (and hence the amount of additional experiments/GPU hours/costs needed). We nevertheless expect that these factors can be incorporated quite naturally into our scaling laws which are power-law based.
Theory. Providing a theoretical explanation of the scaling laws is an important open question in the field, but existing work has been focusing on relatively simple theoretical setups like power-law random feature models to reproduce power-law scaling in deep neural networks maloney2022solvable ; bordelon2024dynamical ; paquette20244+ . To our best knowledge, there is no existing theory that can explain the change of scaling exponents with respect to the amount of pretraining data. In the current work, we have focused on the practical aspects of the scaling laws, following previous influential scaling law work kaplan2020scaling ; hoffmann2022training , and have not attempted to provide a theoretical explanation of the scaling laws as the current theoretical understanding of scaling laws is still limited. Nevertheless, it would be interesting to study if such a change of scaling exponents can be explained by existing theories or if new theoretical frameworks are needed.

### v2

@@ 3.1 Functional Forms for the Scaling Laws
To recite, our goal is to find a functional form $L(D_{1},D_{2})$ that captures how the second-stage loss depends jointly on both stages. To derive a principled formulation, we impose two natural constraints:
@@
$L(D_{1},D_{2})=L_{D_{1}}(D_{2})=A_{D_{1}}D_{2}^{-\alpha_{D_{1}}}+E_{D_{1}}.$ (4)
This is consistent with classical neural scaling laws and Observation 1. It reflects the expectation that, when starting from a fixed initialization, additional tokens improve performance predictably.
Condition 2: For any fixed value of $D_{2}$, the loss should exhibit power-law behavior with respect to the number of first-stage tokens $D_{1}$:
$L(D_{1},D_{2})=L_{D_{2}}(D_{1})=A_{D_{2}}D_{1}^{-\alpha_{D_{2}}}+E_{D_{2}}$ (5)
This is consistent with the power-law ansatz and captures the intuition that a better-trained base model (i.e., larger $D_{1}$) should result in lower loss.
Moreover, this condition implies that as $D_{2}\to 0$, the loss should continuously approach that of the base model: $\lim_{D_{2}\to 0}L(D_{1},D_{2})=AD_{1}^{-\alpha}+E$ (as in Footnote 1, we omit the +1+1 term in $D+1$ expressions for notational simplicity, such that terms proportional to $D_{2}$ are finite as $D_{2}\to 0$), which is a natural requirement for function-preserving model growth; when model capacity is expanded but initialized carefully, the initial loss should remain close to the base model’s loss. We further validate this Condition by showing empirically that the loss follows a power law with respect to $D_{1}$ for fixed $D_{2}$ in Appendix D.
@@
$\textbf{Multiplicative:}\;L(D_{1},D_{2})$ $=AD_{1}^{-\alpha_{1}}D_{2}^{-\alpha_{2}+\alpha_{3}\log D_{1}}+E.$ (6)
$\textbf{Additive:}\;L(D_{1},D_{2})$ $=AD_{1}^{-\alpha_{1}}+FD_{2}^{-\alpha_{2}}+E.$ (7)
$\textbf{Hybrid:}\;L(D_{1},D_{2})$ $=\big(AD_{1}^{-\alpha_{1}}+F\big)D_{2}^{-\alpha_{2}}+E.$ (8)
@@
Note that since $D_{1}^{-\alpha_{1}}D_{2}^{-\alpha_{2}+\alpha_{3}\log D_{1}}=D_{1}^{-\alpha_{2}+\alpha_{3}\log D_{2}}D_{2}^{-\alpha_{2}}$, the multiplicative form with an interaction term satisfies our conditions, while other forms do not have such an interaction term. We further note that multiplicative scaling laws have been observed in Mikami et al. (2022); Zhang et al. (2024a) but without the interaction term. Liew et al. (2025) empirically showed that sparse upcycling (training sparse mixture-of-experts (MoE) models reusing existing dense models) follows a scaling law similar to Equation 6 but under different motivation and conditions. Our work directly extends their findings to other pretraining paradigms. Equation 8 has also been studied in Barnett (2024).
3.2 Fitting Scaling Laws
To determine the functional form described in the previous Section that best describes the scaling behavior of model reuse methods, we consider a variety of datasets and methods beyond those described in the previous Section.
CPT. We consider both code and mathematics datasets and perform CPT on them as in the previous Section. Additionally, we consider the following variants of CPT that are commonly used in practice (Ibrahim et al., 2024) and test them on the code dataset:
@@ 3.2 Fitting Scaling Laws
  • CPT with replay (rep): A portion of the first-stage data is mixed into the second-stage data, which is common for mitigating catastrophic forgetting. We consider a replay ratio of 0.25, i.e., 25% of the second-stage data is from the first stage.
@@
  • CPT from stable phase (sta): We consider CPT starting from a base model checkpoint in the stable phase of the WSD learning rate schedule. This is to avoid adverse effects from re-warming the learning rate from a decayed value.
Model growth. In addition to depth-wise stacking described in the previous Section, we consider width expansion that double the size of the base model. For a growth factor of 2, we increase the size of the hidden dimension by $\sqrt{2}$ for width expansion. We further consider stacking variants:
  • Larger growth factor (x4): We consider a larger growth factor of 4 times (instead of 2).
@@
  • Model growth from stable phase (sta): Similar to CPT, we train the stacked model starting from a base model checkpoint in the stable phase.
Comparing various functional forms. Running the experiments as described above and obtaining the results, we fit the losses with the functional forms of Equations 6, 7, and 8. Following Hoffmann et al. (2022); Besiroglu et al. (2024), we perform optimization using the Huber loss ($\delta=10^{-3}$) and the BFGS algorithm, to fit the logarithm of the loss via the LogSumExp trick applied to the RHS of functional forms. The leave-one-out root mean square error (RMS) serves as the goodness-of-fit metric.
As can be seen in Table 1, the multiplicative form with interaction consistently achieves the lowest error across all methods and datasets, indicating that it best captures the underlying scaling behavior. To visualize these relationships, we compare the loss-versus-token plots of our proposed scaling law against alternative functional forms (Figure 3, left panel), using a 0.5B-to-1B model by stacking. The predicted losses from the alternative forms are visually confirmed to fit the empirical data worse. More importantly, the predicted losses from alternative forms with varying $D_{1}$ remain nearly parallel across the range of $D_{2}$, whereas ours converges, a distinct behavior that is not captured by other forms. In Appendix D.2, we compare with yet another scaling formulation (“continuous” scaling law (Que et al., 2024; Tissue et al., 2024)) and show that our scaling law still fits better.
Table 1: Multiplicative scaling law with interaction consistently achieves lowest error. Leave-one-out RMS error ($\times 10^{-3}$) for fitting the loss for model reuse of a 0.1B base model. Functional forms are from Equations 6 (and the specific case where $\alpha_{3}=0$), 7, and 8.
RMS $(\times 10^{-3})$ Code Math CPT (rep) CPT (sta) Exp x2 Stk x2 Stk x4 Stk (sta)
Mul. 1.573 1.737 0.987 1.371 2.790 2.845 2.152 2.957
Mul. ($\alpha_{3}=0$) 2.095 3.147 2.222 2.366 4.837 7.818 6.557 8.757
Add. 3.913 7.443 4.646 4.315 9.855 10.323 8.866 10.559
Hyb. 2.245 3.340 2.223 2.367 4.841 7.748 6.667 8.822
Figure 3: Left: Visual comparison of various functional forms. We show scaling behavior of second-stage training tokens ($D_{2}$) for different values of first-stage tokens ($D_{1}$) for model growth by stacking (growth factor 2, 0.5B-to-1B model), comparing various functional forms. Our proposed multiplicative scaling law with interaction (solid lines) fits the data (dots) best, and lines with different $D_{1}$ cross over, indicating saturation effects. Note that the hybrid functional form (dotted lines) nearly overlaps with the multiplicative without interaction form (dashed lines) in this plot. Middle: Joint scaling law fit for continual pretraining on code data. Orange points indicate the 10% of data with lowest losses used for validation. Right: Stacking with extremely large token-to-parameter ratio. At the limit of large token ($D_{2}$)-to-parameter ratio, the values of the loss are too noisy to yield a definite characterization of the crossover region.
Overall, the proposed scaling law is robust across datasets, configurations, and methods investigated. While we acknowledge the possibility of edge cases that deviate from this law, it is important to note that our empirical validation specifically targeted training methodologies that are considered scalable and representative of current best practices in the field. Consequently, we conclude that the scaling behavior and saturation effect identified here are not isolated phenomena, but are fundamentally representative of general model reuse methods employed in efficient LLM development.
4 Extensions and Limitations
@@ 4.1 Joint Scaling Incorporating Model Size
We proceed to extend the data scaling law to include model size, $N$. Let us first focus on CPT where the model size remains unchanged after second-stage pretraining. To determine the functional form, we impose that the loss follows the well-grounded “Chinchilla” scaling law (Hoffmann et al., 2022) (which jointly models the base model’s loss with respect to dataset and model sizes) with respect to $D_{2}$ and $N$.
Condition 3: For a fixed base model trained on $D_{1}$ tokens, the second-stage loss should follow the Chinchilla scaling law with respect to the number of tokens $D_{2}$ and model size $N$:
$L_{D_{1}}(D_{2},N)=A_{D_{1}}D_{2}^{-\alpha_{D_{1}}}+N_{D_{1}}^{-\beta_{D_{1}}}+E_{D_{1}}.$
@@
Fitting the joint scaling law. We conduct experiments with base models of sizes 15 million (M), 44M, 0.1B, 0.2B, 0.5B and 1B, training them with different numbers of first and second-stage tokens as in Section 3.2. To fit the joint scaling law, we use the same fitting procedure as in Section 3.2, additionally taking $N$ as an additional variable. In the left panel of Figure 4, we show that the formula fits the data well for CPT on code. See Table 7 in Appendix D.3 for the fitted coefficients. In the same Appendix, we further extend our results to model growth where model sizes increase.
The joint scaling law allows us to extrapolate performance to models and datasets larger than those studied experimentally (Figure 3, middle panel). We show in Appendix E.1 with a specific example how it can be used to predict the performance of larger models and datasets. We can further study the compute optimality of CPT and model growth using the joint scaling law, as shown in Appendix E.2.
Figure 4: Left: Overtrained models have larger gradient norms when undergoing model reuse. We show the gradient norm curves for continual pretraining on code data (1B model) with respect to number of training steps in the second stage for base models trained with different values of first-stage tokens ($D_{1}$). We see that overtrained models have larger gradient norms, indicating that the optimization difficulty increases with $D_{1}$, which may explain the saturation effect observed in the scaling law. See Appendix E.4 for model growth results, where similar trends are observed. Right: Loss versus second-stage (from-scratch) training token for stacking (training from scratch) a 0.5B-to-1B model (1B model). Data points and fitted lines using our multiplicative scaling law with different number of first-stage training tokens (power-law) are shown. It can be seen that as the number of training tokens increases, the losses from model growth saturate, and from-scratch training eventually outperforms it.
4.2 Limitations of the Scaling Law
We dicusss scenarios where our scaling laws may not hold, especially when the number of training tokens is extremely large.
Large $D_{1}$ limit. A theoretical limitation of Equation 1 is that as $D_{1}$ increases, the effective scaling exponent for $D_{2}$, $\alpha_{\text{eff}}=\alpha_{2}-\alpha_{3}\log~D_{1}$, decreases without bound, potentially becoming negative when $D_{1}>e^{\alpha_{2}/\alpha_{3}}$. However, based on our fitted coefficients where $\alpha_{2}$ often exceeds $\alpha_{3}$ by at least an order of magnitude (see Table 7 in Appendix D.3), this theoretical threshold (e.g., $D_{1}>10^{17}$) is unlikely to be reached in practice. We anticipate the scaling law would break down well before this point.
Alternatively, one can formally model this inevitable thresholding behavior, which marks a deviation from the core power-law ansatz. We propose a modification motivated by Clark et al. (2022) (used for MoE scaling laws): In Equation 1, we replace $D_{1}$ with a saturated form, $\hat{D_{1}}$, defined as
$\hat{D_{1}}=\left(D_{1}^{-1}+D_{\rm max}^{-1}\right)^{-1},$ (10)
where $D_{\rm max}$ ($\lesssim e^{\alpha_{2}/\alpha_{3}}$) is a constant representing the maximum effective number of tokens that can be learned from the first stage. This modification is designed such that when $D_{1}\gg D_{\rm max}$, $\hat{D_{1}}$ saturates to $D_{\rm max}$, preventing unbounded growth of the effective exponent.
Large $D_{2}$ limit. Our scaling laws predict that base models trained with different number of tokens in the first stage ($D_{1}$) will converge at some point, and diverge again afterwards, with smaller $D_{1}$ models performing better at large $D_{2}$. To see if the latter occurs (under reasonable compute constraints), we train a 15M model with different $D_{1}$ values, stack them with growth factor 2, and continue training with extremely large number of tokens in terms of token-to-parameter ratio (up to around a factor of 2000); see Figure 3, right panel.
However, we find that we effectively enter a noise floor at such limits, where stochastic noise can dominate the results. To disentagle such effects, we decay the checkpoints multiple times with different random seeds to estimate the noise variance (details in Appendix E.3). Despite these measures, the data remained too noisy to yield a definitive conclusion regarding the post-crossover regime. We also do not observe crossover for a larger 0.5B-to-1B model; see Appendix E.3. In summary, while our scaling laws predict a crossover, we do not yet find clear empirical evidence that they hold once this point is passed. Given the noise and the exponentially larger compute required to explore this territory, we leave this significantly more compute-intensive study as an open problem.
5 Towards Understanding and Mitigating the Saturation Effects
5.1 Mechanistic Causes
To understand the optimization dynamics of overtrained models, we analyze gradient norms during the second stage of pretraining. We show in the left panel of Figure 4 the gradient norm curves of overtrained and well-trained models during model reuse. We find that overtrained models exhibit larger gradient norms generally, providing a form of mechanistic explanation for the saturation phenomenon. This aligns with some previous findings (Klein et al., 2024), and also aligns with our initial intuition that the over-optimized parameters are situated in a highly steep/sensitive part of the parameter space relative to the second-stage training, making further optimization difficult. We have also investigated other possible observables but do not find strong correlation with overtraining; see Appendix E.4.
5.2 Practical Mitigation Strategy
Regularization techniques like gradient clipping and layer normalization, are effective at controlling gradient norms, which could potentially mitigate saturation effects. However, such techniques are already employed in our models. Instead of pursuing further optimization or architecture modifications, which may be impractical due to the complexity of large-scale training of language models, we adopt a direct, data-driven approach: leveraging the derived scaling laws for both second-stage and from-scratch pretraining to guide practitioners in minimizing the negative effects of overtraining. We specifically target the stacking-based model growth scenario, as it exhibits more serious saturation effects and training from scratch is a common practical alternative.
Denote $L^{\rm scratch}_{2N}(D)$ by the loss of a model of size $2N$ trained from scratch for $D$ tokens, and $L^{\rm growth}_{N}(D_{1},D_{2})$ by the loss of a model of size $2N$ grown from a base model of size $N$ trained for $D_{1}$ tokens, and then trained for $D_{2}$ tokens.
In Figure 4 (left panel), we show the plots of the two losses, $L^{\rm growth}_{N}(D_{1},D_{2})$, $L^{\rm scratch}_{2N}(D)$ with $D=D_{2}$, using our experimental data as well as the fitted scaling laws for 0.5B-to-1B grown model and from-scratch 1B model training. We see that for small $D$ values, model growth outperforms from-scratch training for all $D_{1}$ values considered. However, as $D$ increases, the losses from model growth saturate, and from-scratch training eventually outperforms it. Consequently, the optimal choice between model growth and from-scratch training depends directly on the initial investment ($D_{1}$) and the total target training tokens ($D_{2}$ or $D$). This analysis provides a simple, scaling-law-based strategy for practitioners to avoid the negative consequences of reusing overtrained base models.
5.3 A Toy Model
Our previous results are mainly observational and we are unaware of existing theories (Bahri et al., 2024; Paquette et al., 2024; Maloney et al., 2022) that can directly explain why the second-stage scaling exponent decreases logarithmically with $D_{1}$. We henceforth take an alternative view by proposing a minimal toy setup in which the logarithmic saturation of the scaling exponent emerges, drawing on the work of learning curve theory (Hutter, 2021), maximum entropy principle (Mandelbrot and others, 1953; Jaynes, 1957b, a), and power-law assumptions. Detailed derivations are relegated to Appendix E.5.
We model the feature space (of infinite dimension) as a $k$-ary tree with branching factor $k$. Each datum consists of a single feature, and correct learning of a feature requires observing it once (Hutter, 2021). The rank $r$ of a feature at depth $d$ is approximately $r\approx k^{d}$. We assume the cost of learning a feature at depth $d$ is proportional to $d\approx\frac{\log r}{\log k}$. Then, following maximum entropy principle, the probability of finding a feature at depth $d$ decays exponentially ($\lambda>0$ is a constant):
$P(d)\propto e^{-\lambda d}\propto r^{-\left(\frac{\lambda}{\log k}\right)},$
establishing that the power-law exponent is $\alpha=\frac{\lambda}{\log k}$.
Hutter (2021) showed that if the data distribution follows power law with exponent $\alpha+1$, then the overall learning curve also follows a power law with exponent $\alpha$ for small $\alpha$’s. This sets the exponent of learning with $D_{2}$. To model the effect of $D_{1}$, we posit that first-stage pretraining uncovers more features, effectively expanding the effective branching factor $k$ by a power law: $k_{new}\approx k\cdot D_{1}^{\delta}$, where $\delta>0$ is a constant. Then, the new exponent becomes
$\alpha_{new}\approx\frac{\lambda}{\log k+\delta\log D_{1}}\approx\alpha-\theta\log D_{1},$
with $\theta=\frac{\lambda\delta}{\left(\log k\right)^{2}}$. We have thus shown (in a minimal setup) that saturation in the scaling exponent can emerge from a tree model with its branching factor modified by first-stage training.
6 Conclusion
We have provided a fairly broad scaling study of two-stage pretraining and there are other potential future directions worth diving deeper into, besides the ones we preliminarily studied in the main text (post-crossover region, mechanistic explanation, theory for saturated scaling): (1) Incorporating other factors into the scaling laws, e.g., replay ratio (Que et al., 2024; Wang et al., 2025) and the growth factor in model growth (Du et al., 2024). (2) Developing more advanced and scalable techniques to completely eliminate saturation effects, complementing the scaling-law-based guidelines presented in this paper.
Impact Statement
This paper works toward the goal of advancing the field of machine learning and language modeling, with an emphasis on scaling. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.